### v2

@@ IV.3 Two-body scattering of $D^{(*)}\bar{B}^{(*)}$ systems
In addition to the binding energies, we further derive information about $D^{(*)}\bar{B}^{(*)}$ elastic scattering by solving the Lippmann-Schwinger equation. In quantum theory, the scattering rate represents the probability per unit time for a scattering event, it is proportional to the cross section, the number of targets, and the flux Thomson:2013zua . Using Eqs. (12)-(16), we now calculate the scattering rates between $D^{(*)}$ and $\bar{B}^{(*)}$. The variations of the scattering rates $k\sigma(k)$ with cutoffs are shown in Figs. 9 and 10, as functions of the center-of-mass energy.
@@
Let us focus on the scattering rates $k\sigma(k)$ for the $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ systems in Figs. 9(a) and 9(b). The enhancement of $k\sigma(k)$ near the threshold is gradual, which is insufficient to confirm the presence of a bound or virtual state. Look at the scattering lengths and effective ranges in Table 12, We find that the scattering lengths $a$ for different cutoffs are negative, which indicates that the phase shifts satisfy $20\leq\delta<\pi/$ 2 Zhai:2021uap . As a result, there are no evident bound states in the two systems, which aligns with the outcomes found in the previous subsection by solving the Schrödinger equation.
@@
In Table 12, we find that, in the $D^{*}\bar{B}^{*}[0(0^{+})]$ system, the scattering lengths are positive for different cutoff values, indicating the presence of unambiguous bound states, which supports the calculations in above subsection.
@@
Figure 10: The dependence of the $S$-wave scattering rate $k\sigma$ for the $D^{*}\bar{B}^{*}[0(0^{+})]$, $D^{*}\bar{B}^{*}[0(1^{+})]$, $D^{*}\bar{B}^{*}[0(2^{+})]$ systems on the cutoff $\Lambda$.
For the $D^{*}\bar{B}^{*}[0(2^{+})]$ system at different cutoffs, the effective interaction is attractive at small momentum (large distance) but repulsive at large momentum (small distance). It initially causes a positive phase shift at low momentum and then a negative phase shift at high momentum, so that the scattering rate drops to zero, and then gradually increases against the center-of-momentum energy. The negative scattering length indicates that the attractive effect is not strong enough to bind $D^{*}$ and $\bar{B}^{*}$ together in $0(2^{+})$ channel.
Table 12: The scattering length $a$ and effective range $r$ in the five $D^{(*)}\bar{B}^{(*)}[I=0]$ scattering channels. The cutoff parameter $\Lambda$, $a$, and $r$ are in units of $\rm{GeV}$, $\rm{fm}$, and $\rm{fm}$ respectively.
@@ IV.4 Estimation of the contact term LECs
Recently, a lattice QCD study conducted by Alexandrou et al. Alexandrou:2023cqg found shallow bound states for both $J=0$ and $J=1$ $\bar{b}\bar{c}ud$ tetraquarks, with binding energies of $0.5_{-1.5}^{+0.4}\rm{MeV}$ and $2.4^{+2.0}_{-0.7}~{}\rm{MeV}$, respectively. Padmanath et al. reported a lattice QCD research indicating that an $I(J^{P})=0(1^{+})$ $\bar{b}\bar{c}ud$ bound state with a binding energy of $43(^{+7}_{-6})(^{+24}_{-14})~{}\rm{MeV}$ can exist below the $B^{*}\bar{D}$ threshold Padmanath:2023rdu . And, they presented another study searching for tetraquark candidates with exotic quark content $bc\bar{u}\bar{d}$ in the $I=0$ and $J^{P}=0^{+}$ channel, and found a sub-threshold pole in the $S$-wave $D\bar{B}$ scattering amplitude, which corresponds to the binding energy of $39(^{+4}_{-6})(^{+8}_{-18})\rm{MeV}$ below the $D\bar{B}$ threshold Radhakrishnan:2024ihu . The above results provide valuable references for redetermining the contact terms’ LECs.
To ascertain the possible parameter region which allows bound states exist in the $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ systems at $\Lambda=0.5~{}\rm{GeV}$, we can vary the LECs $D_{a}$ and $E_{a}$ within the ranges $[-80,10]~{}\rm{GeV^{-2}}$ and $[-50,-10]~{}\rm{GeV^{-2}}$, respectively. We find that when the LECs satisfy the relation:
@@
a bound state with a binding energy of 30 $\rm MeV$ below the $D\bar{B}$ threshold can be obtained. The green band in Fig. 11 represents the corresponding parameter region. Similarly, we can also determine the contact term LECs for the $D\bar{B}^{*}[0(1^{+})]$ system. For binding energy of $0~{}\rm{MeV}$, we have:
@@
Figure 11: The binding energies of the $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ states vary with the redetermined contact LECs $D_{a}$ and $E_{a}$ with cutoff $\Lambda=0.5~{}\rm{GeV}$. The purple and green bands correspond to the $D\bar{B}$ and $D\bar{B}^{*}$ systems, respectively. The black solid line represents the contact LECs determined by the resonance saturation model. The parallel dashed (dotted) lines at the boundaries of the purple (green) band stand for the regions of parameters within the binding energies $0~{}\rm{MeV}$ and $30~{}\rm{MeV}$. The arrow stands for the direction that binding energy increases for both $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ systems.
@@
In this work, we systematically investigate the interactions of the $S$-wave $D^{(*)}\bar{B}^{(*)}$ systems within the framework of ChEFT using the heavy hadron formalism. We calculate the effective potentials, including contributions from the contact, OPE, and TPE terms, up to NLO at the one-loop level, adopting Weinberg’s formalism. By performing a Fourier transformation, we analyze the behaviors of the effective potentials in coordinate space in detail and employ the Gauss form factor to regularize the divergence in the integral at large momentum. We then insert the coordinate space potentials into the Schrödinger equation to search for potential $D^{(*)}\bar{B}^{(*)}$ bound states.
@@
The results show that all total potentials in the $I=1$ channels are repulsive, whereas those in the $I=0$ channels are attractive. Further calculations examine the variation of interactions with the cutoff parameter $\Lambda$ across five $I=0$ $D^{(*)}\bar{B}^{(*)}$ channels. We find that bound states are more likely to exist in the $D\bar{B}[0(0^{+})]$, $D^{*}\bar{B}^{*}[0(0^{+})]$, and $D^{*}\bar{B}^{*}[0(1^{+})]$ channels. Notably, the TPE contributions are more sensitive to the cutoff in the $I=0$ $D^{*}\bar{B}^{*}$ channels and play a dominant role in these channels, unlike the case in $D\bar{B}[0(0^{+})]$. We list the mass and binding energies in Table 9.
@@
We also discuss the $S$-wave coupled-channel effects of $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ and $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ systems. The numerical results listed in Table 10 indicate that, for $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ system, $D\bar{B}[0(0^{+})]$ channel is dominant, and the coupled channels can somewhat strengthen the binding of the single channel. For $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ system, $D\bar{B}^{*}[0(1^{+})]$ is dominant, and we obtain no binding solution, which is consistent with the single-channel calculation. In a word, the inclusion of the coupled channels does not change our main conclusions given in the single-channel case.
@@
By substituting the momentum space potentials into the Lippmann-Schwinger equation, we calculate the scattering $T$-matrices and the associated scattering phase shifts. To gain further insight into the $S$-wave $D^{(*)}\bar{B}^{(*)}$ interactions, we calculate the scattering rate, scattering length, and effective range, and provide their numerical results in Table 12 based on the $T$-matrix and scattering phase. Accordingly, we study the dependence of these physical quantities on the cutoff parameter $\Lambda$. In our investigation, the shallow bound state is more likely to exist in the $D\bar{B}[I(J^{P})=0(0^{+})]$ system than in the $D\bar{B}^{*}[I(J^{P})=0(1^{+})]$ system.
@@
Based on the above calculations, we find that the interactions in the $D^{*}\bar{B}^{*}[I=0]$ system are more attractive than those in the $D\bar{B}^{(*)}[I=0]$ systems, and $D^{*}\bar{B}^{*}[I(J^{P})=0(0^{+})]$ and $D^{*}\bar{B}^{*}[I(J^{P})=0(1^{+})]$ systems possess large binding energies and positive scattering lengths, which suggests strong bound state formations in these channels. Therefore, we strongly recommend the experiment to find the $D^{*}\bar{B}^{*}[I=0]$ tetraquark systems.
@@
Considering other theoretical and lattice QCD studies, we redetermined the contact term LECs and calculated the effective potentials for the 0($0(0^{+})$ $D\bar{B}$ system and the $0(1^{+})$ $D\bar{B}^{*}$ system. In Fig. 11, we present the dependencies of the binding energies of these two charm-bottom systems on the redetermined contact LECs $D_{a}$ and $E_{a}$.
@@ Acknowledgments
Zhe Liu expresses sincere gratitude to Prof. Zhan-Wei Liu for his insightful discussions and extends thanks to Zi-Le Zhang and Ri-Qing Qian for their valuable contributions to the discussions. This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335001, 12247101, 12465016 and 12005168, the “111 Center” under Grant No. B20063, the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 22JR5RA171), the fundamental Research Funds for the Central Universities (Grant No. lzujbky-2023-stlt01), and the project for top-notch innovative talents of Gansu province.
@@ Appendix A Coupled-channel calculations
In this section, we show the coupled-channel calculations of $D^{(*)}\bar{B}^{(*)}$ systems. The effective potentials derived from Lagrangians in Eq. 1 and Eq. 6 for $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{(*)}\bar{B}^{(*)}\to D^{*}\bar{B}^{*}$ scattering are presented in this sections.
@@ A.1 $D\bar{B}\to D^{*}\bar{B}^{*}$
Figure 12: LO contact, NLO contact, LO OPE, NLO OPE and NLO TPE diagrams of the process $D\bar{B}\to D^{*}\bar{B}^{*}$.
@@
The one-loop corrections to the contact amplitude are listed as:
@@
×$\times(\epsilon^{*}_{3}\cdot\epsilon^{*}_{4}),$ (117)
@@
×$\times(\epsilon^{*}_{3}\cdot\epsilon^{*}_{4}),$ (119)
@@
where coefficients appearing in the above contact amplitudes are shown in Table 13. The NLO OPE diagrams are illustrated in Fig. 12, and the corresponding amplitudes:
@@
$\mathcal{M}_{b4.3}^{(2)}=$ $-4\frac{g_{1}^{3}g_{2}}{f^{2}}AJ_{22}^{g}\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (123)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (127)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (128)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (129)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (130)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (131)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (132)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (133)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (134)
@@
×$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4}.$ (135)
The TPE amplitudes at NLO are given by:
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{a4.1(2)}$ 00 00 $-\delta_{2}$ $0(\delta_{2})$
$A_{a4.3(4)}$ 00 00 $-\delta_{1}$ $0(\delta_{1})$
@@
$A_{a4.6}$ $\frac{1}{4}(D_{a}+E_{a}$ $+2D_{b}+2E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a}$ $+2D_{b}-6E_{b})$ $-\delta_{1}$ $-\delta_{2}$
$A_{a4.7}$ 00 00 00 00
$A_{a4.(10+11)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{1}$
$A_{a4.12}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $-\delta_{1}$ 00
$A_{a4.(13+14)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{2}$ 00
$A_{a4.15}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $-\delta_{2}$
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{b4.1(2)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $-\delta_{2}$ $0(\delta_{2})$
$A_{b4.3(4)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $-\delta_{1}$ $0(\delta_{1})$
$A_{b4.5(6)}$ $-\frac{1}{12}$ $\frac{1}{4}$ 00 00
$A_{b4.7}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b4.(8+9)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{1}$
$A_{b4.10}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\delta_{1}$ 00
$A_{b4.(11+12)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{2}$
$A_{b4.13}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\delta_{2}$ 00
$A_{b4.14(15)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b4.16(17)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00

### v3

@@ IV.3 Two-body scattering of $D^{(*)}\bar{B}^{(*)}$ systems
In addition to the binding energies, we further derive information about $D^{(*)}\bar{B}^{(*)}$ elastic scattering by solving the Lippmann-Schwinger equation. In quantum theory, the scattering rate represents the probability per unit time for a scattering event, and it is proportional to the cross section, the number of targets, and the flux Thomson:2013zua . Using Eqs. (12)-(16), we now calculate the scattering rates between $D^{(*)}$ and $\bar{B}^{(*)}$. The variations of the scattering rates $k\sigma(k)$ with cutoffs are shown in Figs. 9 and 10, as functions of the center-of-mass energy.
@@
Let us focus on the scattering rates $k\sigma(k)$ for the $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ systems in Figs. 9(a) and 9(b). The enhancement of $k\sigma(k)$ near the threshold is gradual, which is insufficient to confirm the presence of a bound or virtual state. Looking at the scattering lengths and effective ranges in Table 12, we find that the scattering lengths $a$ for different cutoffs are negative, which indicates that the phase shifts satisfy $0\leq\delta<\pi/2$ Zhai:2021uap . As a result, there are no evident bound states in the two systems, which aligns with the outcomes found in the previous section by solving the Schrödinger equation.
@@
In Table 12, we find that, in the $D^{*}\bar{B}^{*}[0(0^{+})]$ system, the scattering lengths are positive for different cutoff values, indicating the presence of unambiguous bound states, which supports the calculations in the above section.
@@
Figure 10: The dependence of the $S$-wave scattering rate $k\sigma$ for the $D^{*}\bar{B}^{*}[0(0^{+})]$, $D^{*}\bar{B}^{*}[0(1^{+})]$, and $D^{*}\bar{B}^{*}[0(2^{+})]$ systems on the cutoff $\Lambda$.
For the $D^{*}\bar{B}^{*}[0(2^{+})]$ system at different cutoffs, the effective interaction is attractive at small momentum (large distance) but repulsive at large momentum (small distance). It initially causes a positive phase shift at low momentum and then a negative phase shift at high momentum, so that the scattering rate drops to zero, and then gradually increases against the center-of-momentum energy. The negative scattering length indicates that the attractive effect is not strong enough to bind $D^{*}$ and $\bar{B}^{*}$ together in the $0(2^{+})$ channel.
Table 12: The scattering length $a$ and effective range $r$ in the five $D^{(*)}\bar{B}^{(*)}[I=0]$ scattering channels. The cutoff parameter $\Lambda$, $a$, and $r$ are in units of $\rm{GeV}$, $\rm{fm}$, and $\rm{fm}$, respectively.
@@ IV.4 Estimation of the contact term LECs
Recently, a lattice QCD study conducted by Alexandrou et al. Alexandrou:2023cqg found shallow bound states for both $J=0$ and $J=1$ $\bar{b}\bar{c}ud$ tetraquarks, with binding energies of $0.5_{-1.5}^{+0.4}\rm{MeV}$ and $2.4^{+2.0}_{-0.7}~\rm{MeV}$, respectively. Padmanath et al. reported a lattice QCD study indicating that an $I(J^{P})=0(1^{+})$ $\bar{b}\bar{c}ud$ bound state with a binding energy of $43(^{+7}_{-6})(^{+24}_{-14})~\rm{MeV}$ can exist below the $B^{*}\bar{D}$ threshold Padmanath:2023rdu . And, they presented another study searching for tetraquark candidates with exotic quark content $bc\bar{u}\bar{d}$ in the $I=0$ and $J^{P}=0^{+}$ channels, and found a subthreshold pole in the $S$-wave $D\bar{B}$ scattering amplitude, which corresponds to the binding energy of $39(^{+4}_{-6})(^{+8}_{-18})\rm{MeV}$ below the $D\bar{B}$ threshold Radhakrishnan:2024ihu . The above results provide valuable references for redetermining the contact terms’ LECs.
To ascertain the possible parameter region which allows bound states to exist in the $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ systems at $\Lambda=0.5~\rm{GeV}$, we can vary the LECs $D_{a}$ and $E_{a}$ within the ranges $80,10]~\rm{GeV^{-2}}$ and $50,-10]~\rm{GeV^{-2}}$, respectively. We find that when the LECs satisfy the relation
@@
a bound state with a binding energy of 30 $\rm MeV$ below the $D\bar{B}$ threshold can be obtained. The green band in Fig. 11 represents the corresponding parameter region. Similarly, we can also determine the contact term LECs for the $D\bar{B}^{*}[0(1^{+})]$ system. For binding energy of $0~\rm{MeV}$, we have
@@
Figure 11: The binding energies of the $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ states vary with the redetermined contact LECs $D_{a}$ and $E_{a}$ with cutoff $\Lambda=0.5~\rm{GeV}$. The purple and green bands correspond to the $D\bar{B}$ and $D\bar{B}^{*}$ systems, respectively. The black solid line represents the contact LECs determined by the resonance saturation model. The parallel dashed (dotted) lines at the boundaries of the purple (green) band stand for the regions of parameters within the binding energies $0~\rm{MeV}$ and $30~\rm{MeV}$. The arrow stands for the direction that binding energy increases in for both $D\bar{B}[0(0^{+})]$ and $D\bar{B}^{*}[0(1^{+})]$ systems.
@@
In this work, we systematically investigated the interactions of the $S$-wave $D^{(*)}\bar{B}^{(*)}$ systems within the framework of ChEFT using the heavy hadron formalism. We calculated the effective potentials, including contributions from the contact, OPE, and TPE terms, up to NLO at the one-loop level, adopting Weinberg’s formalism. By performing a Fourier transformation, we analyzed the behaviors of the effective potentials in coordinate space in detail and employed the Gauss form factor to regularize the divergence in the integral at large momentum. We then inserted the coordinate space potentials into the Schrödinger equation to search for potential $D^{(*)}\bar{B}^{(*)}$ bound states.
@@
The results showed that all total potentials in the $I=1$ channels are repulsive, whereas those in the $I=0$ channels are attractive. Further calculations examined the variation of interactions with the cutoff parameter $\Lambda$ across five $I=0$ $D^{(*)}\bar{B}^{(*)}$ channels. We found that bound states are more likely to exist in the $D\bar{B}[0(0^{+})]$, $D^{*}\bar{B}^{*}[0(0^{+})]$, and $D^{*}\bar{B}^{*}[0(1^{+})]$ channels. Notably, the TPE contributions are more sensitive to the cutoff in the $I=0$ $D^{*}\bar{B}^{*}$ channels and play a dominant role in these channels, unlike the case in $D\bar{B}[0(0^{+})]$. We list the mass and binding energies in Table 9.
@@
We also discussed the $S$-wave coupled-channel effects of $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ and $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ systems. The numerical results listed in Table 10 indicate that, for the $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ system, the $D\bar{B}[0(0^{+})]$ channel is dominant, and the coupled channels can somewhat strengthen the binding of the single channel. For $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ system, $D\bar{B}^{*}[0(1^{+})]$ is dominant, and we obtained no binding solution, which is consistent with the single-channel calculation. In a word, the inclusion of the coupled channels does not change our main conclusions given in the single-channel case.
@@
By substituting the momentum space potentials into the Lippmann-Schwinger equation, we calculated the scattering $T$-matrices and the associated scattering phase shifts. To gain further insight into the $S$-wave $D^{(*)}\bar{B}^{(*)}$ interactions, we calculated the scattering rate, scattering length, and effective range, and provide their numerical results in Table 12 based on the $T$-matrix and scattering phase. Accordingly, we studied the dependence of these physical quantities on the cutoff parameter $\Lambda$. In our investigation, the shallow bound state is more likely to exist in the $D\bar{B}[I(J^{P})=0(0^{+})]$ system than in the $D\bar{B}^{*}[I(J^{P})=0(1^{+})]$ system.
@@
Based on the above calculations, we found that the interactions in the $D^{*}\bar{B}^{*}[I=0]$ system are more attractive than those in the $D\bar{B}^{(*)}[I=0]$ systems, and $D^{*}\bar{B}^{*}[I(J^{P})=0(0^{+})]$ and $D^{*}\bar{B}^{*}[I(J^{P})=0(1^{+})]$ systems possess large binding energies and positive scattering lengths, which suggests strong bound state formations in these channels. Therefore, we strongly recommend the experiment to find the $D^{*}\bar{B}^{*}[I=0]$ tetraquark systems.
@@
Considering other theoretical and lattice QCD studies, we redetermined the contact term LECs and calculated the effective potentials for the $0(0^{+})$ $D\bar{B}$ system and the $0(1^{+})$ $D\bar{B}^{*}$ system. In Fig. 11, we present the dependencies of the binding energies of these two charm-bottom systems on the redetermined contact LECs $D_{a}$ and $E_{a}$.
@@ Acknowledgments
Zhe Liu expresses sincere gratitude to Prof. Zhan-Wei Liu for his insightful discussions and extends thanks to Zi-Le Zhang and Ri-Qing Qian for their valuable contributions to the discussions. This work is supported by the National Natural Science Foundation of China under Grant No. 12335001, No. 12247101, No. 12465016, and No. 12005168, the “111 Center” under Grant No. B20063, the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 22JR5RA171), the fundamental Research Funds for the Central Universities (Grant No. lzujbky-2023-stlt01), and the project for top-notch innovative talents of Gansu province.
VI DATA AVAILABILITY
The data that support the findings of this article are openly available Zhe:2025dat .
@@ Appendix A Coupled-channel calculations
In this appendix, we show the coupled-channel calculations of $D^{(*)}\bar{B}^{(*)}$ systems. The effective potentials derived from Lagrangians in Eq. (1) and Eq. (6) for $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{(*)}\bar{B}^{(*)}\to D^{*}\bar{B}^{*}$ scattering are presented in this appendix.
@@ A.1 $D\bar{B}\to D^{*}\bar{B}^{*}$
Figure 12: LO contact ($a4$), LO OPE ($b4$), NLO contact (a4.1a4.15a4.1-a4.15), NLO OPE (b4.1b4.17b4.1-b4.17), and NLO TPE (c4.1c4.2c4.1-c4.2) diagrams of the process $D\bar{B}\to D^{*}\bar{B}^{*}$.
@@
The one-loop corrections to the contact amplitudes are listed as
@@
$\times(\epsilon^{*}_{3}\cdot\epsilon^{*}_{4}),$ (117)
@@
$\times(\epsilon^{*}_{3}\cdot\epsilon^{*}_{4}),$ (119)
@@
where coefficients appearing in the above contact amplitudes are shown in Table 13. The NLO OPE diagrams are illustrated in Fig. 12, and the corresponding amplitudes
@@
$\mathcal{M}_{b4.3}^{(2)}=$ $-4\frac{g_{1}^{3}g_{2}}{f^{2}}AJ_{22}^{g}\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4}$ (123)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (127)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (128)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (129)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (130)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (131)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (132)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (133)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4},$ (134)
@@
$\times\epsilon^{*\mu}_{3}\epsilon^{*\nu}_{4}.$ (135)
The TPE amplitudes at NLO are given by
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{a4.1(2)}$ 0 0 $\delta_{2}$ $0(\delta_{2})$
$A_{a4.3(4)}$ 0 0 $\delta_{1}$ $0(\delta_{1})$
@@
$A_{a4.6}$
$\frac{1}{4}(D_{a}+E_{a}$
$2D_{b}+2E_{b})$
$\frac{-3}{4}(D_{a}-3E_{a}$
$2D_{b}-6E_{b})$
$\delta_{1}$ $\delta_{2}$
$A_{a4.7}$ 0 0 0 0
$A_{a4.(10+11)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{1}$
$A_{a4.12}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{1}$ 0
$A_{a4.(13+14)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{2}$ 0
$A_{a4.15}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{2}$
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{b4.1(2)}$ $\frac{1}{16}$ $\frac{3}{16}$ $\delta_{2}$ $0(\delta_{2})$
$A_{b4.3(4)}$ $\frac{1}{16}$ $\frac{3}{16}$ $\delta_{1}$ $0(\delta_{1})$
$A_{b4.5(6)}$ $\frac{1}{12}$ $\frac{1}{4}$ 0 0
$A_{b4.7}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b4.(8+9)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{1}$
$A_{b4.10}$ $\frac{1}{4}$ $\frac{3}{4}$ $\delta_{1}$ 0
$A_{b4.(11+12)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{2}$
$A_{b4.13}$ $\frac{1}{4}$ $\frac{3}{4}$ $\delta_{2}$ 0
$A_{b4.14(15)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b4.16(17)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0